OpenStudy (liizzyliizz):

Im not sure how to approach this problem. f(1) = 4 and f'(1)= 2 find an equation of the tangent line at x=1 Im not sure how to approach this problem. f(1) = 4 and f'(1)= 2 find an equation of the tangent line at x=1 @Mathematics

OpenStudy (anonymous):

If i had said, "the slope of the line is 2, and the line goes through the point (1,4), what is the equation of the line" could you solve that?

OpenStudy (anonymous):

Use the point slope formula for a line, using f'(1) = 2 as the slope.

OpenStudy (liizzyliizz):

Oh wow.... thats easy then.

OpenStudy (anonymous):

After all the derivative of a function at a point is the slope of the tangent at that point, and the tangent must also pass through the point.

jimthompson5910 (jim_thompson5910):

The equation of the tangent line is y - y1 = m(x - x1) where the tangent line touches the graph of y = f(x) at the point (x1, y1). Also, the slope is m = f'(x1) In your case, (x1, y1) = (1, 4), so x1 = 1 and y1 = 4 Also, m = f'(x1) = f'(1) = 2, so m = 2 Plug all this into y - y1 = m(x - x1) to get y - y1 = m(x - x1) y - 4 = 2(x - 1) Now solve for y y - 4 = 2x - 2 y = 2x - 2 + 4 y = 2x + 2 So the equation of the tangent line is y = 2x + 2

OpenStudy (liizzyliizz):

aha, thank you guys. it was just worded odd to me. *sigh*